OpenStudy (anonymous):

If M(1, 1) is the midpoint of the line segment AB, and if A has coordinates (−2, −4), find the coordinates of B.

OpenStudy (anonymous):

Ok, so you probably know that the midpoint formula is ((x1+x2)/2,(y1+y2)/2)). Assume A is (x1,x2) and you also know that M(1,1) is equal to the resultant x and y values of the midpoint formula. try it and post your answer. I'll check it. Also, if you need more help just ask.

OpenStudy (anonymous):

Yes ; So far i have that 1=(-2+x^2) and 1=(-4+y^2) I am just confused on where to go from there.... ? Assuming that if M(m1,m2) is midpoint of two point being A(x1,y1) and B(x2,y2) then M1=(x1+x2)/2 and M2=(y1+y2)/2 .... i hope that didnt confuse you?

OpenStudy (anonymous):

ok, so you have a good start. However, 1 = (-2+x2)/2 AND 1 = (-4+y2)/2

OpenStudy (anonymous):

yes... but im not sure how to move on past this? i belive i am trying to get the x2 and y2 on the other side of the = sign alone correct?

OpenStudy (anonymous):

yes. i'll solve the x2 and you can solve the y2 part 1= (-2+x2)/2 1*2 = (-2+x2) 2 = (-2+x2) 2+2 = x2 x2 = 4 // i just flipped x2 to the left side for aesthetics.

OpenStudy (anonymous):

y2=6 ??

OpenStudy (anonymous):

that is correct. nice job.

OpenStudy (anonymous):

THANK YOU :) life saver lol :) i totally had like a brain fart but once i saw what you did i knew exactly what i was missing:)

OpenStudy (anonymous):

nice to hear.